HUBBLE SPACE TELESCOPE DAILY REPORT #2206 PERIOD COVERED: 0000Z (UTC) 09/10/98 - 0000Z (UTC) 09/11/98 Daily Status Report as of 254/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/2 7834 (A Search for Giant Planets Around Very Young Nearby Late-type Dwarfs) The Near Infrared Camera (Camera Two) was used to make observations of the very young, late-type, dwarf star HD167605 as part of a search for large planets. Very young late-type dwarfs at distances closer than 30 pc are excellent sites to provide a direct detection of extrasolar planets. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Eleven Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.3 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/1/2/3 7247 (Environments of T-Tauri Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class star HL-Tau. T-Tauri stars are low {0.5-2 solar} mass, pre-main-sequence stars. Many may have circumstellar disks - theoretical models involving accretion disks have successfully explained some of the basic observational properties of these young objects. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.7 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) FGS 7968 () The Near Infrared Camera (Camera One) was used to make imaging observations of the dusty CfA class Seyfert galaxy NGC-3982. These data will allow us to determine if the Seyfert 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/2 7441 (A Search for Zodiacal Dust around Bright Nearby Stars) The Near Infrared Camera (Camera Two) was used to make observations of the zodiacal lights around the bright nearby star Procyon. Dust particles around a main sequence star are quickly removed by Poynting-Robertson drag, therefore, the existence of such dust is an almost certain signature of a current presence of a population of larger bodies in orbit around the star. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/1/2 7858 (The Composition of Kuiper Belt Objects) The Near Infrared Camera (Cameras One and Two) was used to investigate the small body A1996Q66 recently discovered in the outer solar system that appears to be unique both in composition and diversity compared to other classes of solar system objects. The proposal completed as planned, and no anomalies were noted. 1.11 Completed NIC/1/2/3 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera One, Two and Three) was used to make a search for low mass/sub-luminous companions to the low mass M class star PS176. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. As described in HSTAR 6659 and 2.1, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.12 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed Two Sets of NIC/1/2/3 7875 (Searching for Old Stellar Populations in Candidate Proto-Spheroidals) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations searching for old stellar populations in candidate proto-spheroids SA68-5935 and SA68-3307. We have compelling evidence from WF/PC-2 images that these galaxies have optical colors gradients. The near-infrared is necessary, however, to reveal color gradients large enough to separate young and old stellar populations. Only NICMOS imaging provides enough spatial resolution and depth to resolve these gradients. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/1/2/3 7855 (The Evolution of Powerful Radio Galaxies) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the powerful radio galaxy 3C16. Our proposed NICMOS observations combined with our multi-wavelength observing program will finally provide us with the data we need to build a picture of how powerful radio galaxies evolve. The observations were completed as planned, and no problems were reported. 1.15 Completed Two Sets of NIC/2 6767 (A Search for Brown Dwarfs and Luminous Young Planets in the Hyades) The Near Infrared Camera (Camera Two) was used to make observations of the brown dwarf stars RHY166 and RHY126 which are located in the Hyades Cluster. This is to search for substellar companions orbiting low-mass Hyades M class dwarf stars. Differential imaging was used using methane and water band (on/off) filter pairs. The observations were completed as planned, and no problems were reported. 1.16 Completed Two Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.17 Completed NIC/1/2/3 7218 (Nearby Starburst Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the nearby starbursting galaxy NGC-253. This proposal encompasses a study of both the stars and interstellar material in starburst nuclei which will test ideas about the triggering of starbursts and whether starbursts propagate outwards in a nucleus. The observations were completed as planned, and no anomalies were reported. 1.18 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 10 HSTAR 6659 describes the acquisition at 253/112248Z which defaulted to fine lock backup on FGS-1 only when the scan step limit was exceeded on FGS-3. The proposal mentioned in 1.11 may have been affected. The acquisition at 254/011552Z failed when the search radius limit was exceeded on both FGS-1 and FGS-3. The prior FHST full mode update had been good and a FHST map taken at 254/0114928Z showed only slightly higher errors in V3. HSTAR 6661 was written. Three proposals which occur during the period of the next report, NIC 7226, STIS 8062, and WF/PC-2 7909, were affected. Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 22 Successful : 22 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. NICMOS EDAC single bit errors occurred at 253/214356Z and again at approximately 253/2148Z. The NICMOS internal log was dumped at 253/220028Z, when the PRS limit was also adjusted to 12. HSTAR 6660 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH